# DoPTA: Improving Document Layout Analysis using Patch-Text Alignment

### v1

@@ 1 Introduction
The aforementioned reasons highlight the necessity of multimodal modeling for effective document understanding. Currently, transformer architecture has evolved as a ubiquitous framework capable of modeling multiple modalities and has naturally been applied to VDU as well. Most works [46, 47, 17, 1, 41] use a unified transformer approach, wherein image, text, and layout information are processed by a single multi-modal transformer, which is pre-trained with a variety of objectives. The unified transformer based methods focus more on textual information, and treat visual information as secondary. They require extraction of text from a document image using standard OCR techniques, which is later modeled by the unified transformer for downstream tasks. This 2-stage paradigm has two key issues - inflexibility of the OCR pipelines, and error propagation from OCR extraction to downstream tasks. Methods such as Donut [22] instead model both OCR extraction and subsequent understanding of the document in a single end-to-end approach using an encoder-decoder transformer model.
Although these approaches achieve strong performance on semantic tasks such as document question answering [28] and information extraction [38], they fall behind on visual tasks such as document layout analysis, as their primary focus is on modeling the textual features. In this field, state-of-the-art results have been achieved by DiT [23] and VGT [10]. DiT approaches document image understanding in a self-supervised fashion, wherein masked image patches are reconstructed through a ViT encoder, and matched to the tokens from a pre-trained dVAE. VGT builds upon this, adding a Grid Transformer (GiT) to infuse layout information into the image representations. While these methods achieve strong results, we argue that the semantic information from text in the image can be a strong factor to improve the layout understanding. Inspired by the power of contrastive language-image training in representation learning [34] and fine-grained image-text alignment techniques like FILIP [48], we specially design a patch-text alignment loss for documents, using IoU to guide the model to learn effective representations that are semantically and structurally rich. Our key contributions can be summarized as follows:
@@ 2 Related Work
Document Image Understanding. Visual document understanding(VDU) requires careful design of the objectives, owing to the unique structure of these images. The majority of approaches in this field can be categorized into two sub-categories based on the use/non-use of OCR as an input. i) OCR-Based methods include BiVLDoc [27], LayoutLM [46, 47, 17], DocFormer [1], BROS [16], VL-BERT [39], UDOP [41], VGT [10], TILT [33], and UDOC. These works utilize off-the-shelf OCR methods to parse the text and bounding boxes from a document image. The textual and image features are later combined through early or late fusion, using a joint transformer encoder to produce the final representations. Different variants of objectives such as masked image modeling (MIM), masked language modeling (MLM), and image-language alignment are proposed in these papers. However, these works require OCR as an input during inference. ii) OCR-Free methods such as Donut [22], DiT [23], and StructTextv2 [49] instead aim to learn visual features in the absence of OCR as an input during inference, though it may be utilized as a target during pre-training. Donut uses a transformer encoder-decoder architecture with OCR parsing as its pre-training task. On the other hand, DiT learns image features in the absence of any OCR ground truth, by aligning image patches with learned tokens from a discrete VAE tokenizer. StructTextv2 uses a dual objective of image reconstruction and text prediction of masked-out regions. Our work falls into the second category.
@@ 3 Methodology
We now present DoPTA, a novel pre-training method for learning document image representations with strong semantic and structural understanding. The key component of DoPTA is the introduction of a novel fine-grained image-text contrastive alignment objective for document images. This loss imbibes textual-semantic information into the image representations, leading to better structural understanding through the semantics. Despite the strong performance demonstrated by this loss, as shown in Sec. 5DoPTA also includes an image reconstruction loss to incorporate additional structural information. We begin with a description of our model architecture in Sec. 3.1, and introduce the individual image-text contrastive and image reconstruction losses in Sec. 3.2 and Sec. 3.3 respectively. Refer to Figure 1 for an overview of our model.
@@ 3.1 Model architecture
Image Encoder. The  DoPTA encoder (EI) follows is a Vision Transformer [11]. We reshape an image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ into a sequence of flattened 2D patches $\mathbf{x}_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$, where $(H,W)$ is the resolution of the original image, $C$ is the number of channels, $(P,P)$ is the resolution of each image patch, and $L_{I}=HW/P^{2}$ is the resulting number of patches. We randomly mask out a fraction $M$ of the patches by replacing them a learned [MASK] embedding, which are later reconstructed by the image decoder. Learned positional embeddings are added to each patch before passing as input to the transformer. The per-patch embeddings are aligned with both textual and visual information to ensure a rich representation.
Text Encoder. The text encoder (ET) is a transformer model. We extract the text with corresponding bounding boxes from each document image using an off-the-shelf OCR engine. The entire set of texts is concatenated in reading order and tokenized as a single string. We truncate the tokenized string at a maximum sequence length of $L_{T}$ tokens and add learned positional embeddings before passing through the text encoder model. We use the per-token embeddings at the output layer for aligning vision features.
@@ 3.3 Image Reconstruction Loss
Figure 2: Heatmap visualisation of the normalised dot product similarity of image region embeddings with the text embedding for the token ‘phosphine’ taken from DoPTA-HR model. Additional qualitative results are presented in Appendix B.
@@ 4 Experiments
Next, we present experimental results to show the effectiveness of image features produced by DoPTA on a variety of document tasks, including document image classification (Sec. 4.2), document layout analysis (Sec. 4.3), and text detection (Sec. 4.4). We evaluate both variants of our model (224 and 512 resolution), and achieve state-of-the-art results in multiple tasks, outperforming larger models, while adopting a significantly shorter pre-training schedule.
@@ 4.1 Implementation and Pre-Training
We pretrain DoPTA on the IIT-CDIP [37] dataset. This dataset contains 42M pages of black-and-white document images containing rich text. We extract word-level OCR text and their bounding boxes using EasyOCR [19] pipeline and use random cropping as the image augmentation. Though we carefully ensure the quality of the extracted OCR through filtering, some errors in extracted OCR do persist. In Appendix A, we explore the use of a PDF dataset which circumvents this issue. However, we choose the CDIP dataset for pre-training due to its larger scale, and to maintain parity with the baselines, which pretrain on the same dataset. An important point to note is that our method does not require any OCR input during inference.
We use a mix of padded (aspect ratio preserving) and square-cropped images during training to ensure good downstream performance in all settings. We follow the architectural choices of ’CLIP-ViT-B/16’, with our DOPTA encoder and text encoder being 12-layer transformers with 8 attention heads. The hidden (intermediate) sizes are 768 (3072) for the DOPTA encoder and 512 (2048) for the text encoder. Both models are initialized using the ’CLIP-ViT-B/16’ weights. The context length of the text encoder is set to $512$ by linear interpolating the learnt CLIP positional embeddings. As discussed in Sec. 3.1, we adopt a lightweight 2-layer transformer image decoder each with $8$ attention heads. We train at $224\times 224$ resolution, and refer to this model as DoPTA for the remainder of this paper. We also train a high-resolution variant with image resolution $512\times 512$. Going forward, we refer to the high-resolution model as  DoPTA-HR. Both models use a patch size of 16, a global batch size of $2048$, dropout of $0.1$, and learning rate of $31e-$ 3. The masking ratio $M$ for reconstruction is set to $0.6$. We train DoPTA  and DoPTA-HR for 15 epochs ($\approx$250k steps). This is a significantly shorter pre-training schedule compared to other works like DiT [23], LayoutLMv3 [17], and VGT [10], which are pretrained for 500k steps or more.
@@ 4.2 Document Image Classification
We use the RVL-CDIP benchmark to evaluate the classification performance of DoPTA. The benchmark consists of 400K document images split into 320K train, 40K validation and 40K test images. It consists of $16$ different classes like advertisement, email, form, scientific publication, etc. To perform classification, we obtain a single representation embedding per image by average pooling the patch-wise embeddings and directly applying a linear classification head on top. We evaluate both versions of our  DoPTA encoder by finetuning for 100 epochs on the training set as done in DiT [23]. We use AdamW optimiser with a learning rate of $31e-$ 3, a global batch size of 1024 and perform gradient clipping with a value of $0.1$.
Baselines and Results. We compare and report results of classification accuracy on the test set in Table 1. We consider two categories of methods - i) OCR-based methods which rely on the OCR-identified text in the image as input, and ii) OCR-Free methods which treat document image classification purely in the image domain. All results are taken from the DiT[23], except Donut which we finetune ourselves. Donut [22] utilizes an encoder-decoder architecture for end-to-end OCR extraction. To test the performance of Donut in the image domain, we utilize the image encoder alone and evaluate it using the aforementioned setup. While the OCR-based methods achieve the highest performance in this category, we find that both DoPTA and DoPTA-HR outperform all OCR-free methods with similar input resolutions. It is notable that DoPTA  outperforms even the DiT-L model, despite having $<1/3^{rd}$ the parameters, and a much shorter pre-training schedule (250k steps in our case as compared to 500k steps for DiT-L).
@@ 4.3 Document Layout Analysis
For object detection, we use a Cascade R-CNN [4] as the detection pipeline on top of the backbone models, using the Detectron2 [44] library to evaluate our models. We use the same FPN and data processing setup as DiT [23] and VGT [10], with resolution modifying modules at four different transformer blocks ($3$, $5$, $7$ and $11$) to adapt the single-scale ViT to the multi-scale FPN. Let $d$ be the total number of blocks, the $d/3^{rd}$ block is upsampled by 4×4$\times$4 × using a module with $2$ stride-two $2\times 2$ transposed convolution. For the output of the ${d/2}^{th}$ block, we use a single stride-two $2\times 2$ transposed convolution to upsample by 2×2$\times$2 ×. The output of the ${2d/3}^{th}$ block is utilized without additional operations. Finally, the output of $d^{th}$ block is downsampled by 2×2$\times$2 × with stride-two $2\times 2$ max pooling. All images are cropped with probability $0.5$ to a random rectangular patch which is then resized again such that the shortest side is at least $480$ and at most $800$ pixels while the longest is at most $1,333$ pixels.
@@ 4.3.1 PubLayNet
PubLayNet [51] is a large dataset of 360K images for document layout analysis, created from over one million scientific articles in PubMed Central. It includes labeled images with five layout regions: text, title, list, figure, and table. We finetune both versions of our model on the training split $(335,703)$ and evaluate on the validation split $(11,245)$. We follow the setting of DiT [23] and train for 60K steps with a batch size of $16$ and a learning rate of $44e-$ 4.
Baselines and Results. We report the category-wise and overall mean average precision mAP@IoU$[0.50:0.95]$ of bounding boxes in Table 2. We compare with vision-only input models such as DiT and Donut, as well as vision+OCR input models like VGT, LayoutLMv3. DoPTA-HR achieves 94.9 which is on-par with DiT-L despite being less than ${1/3}^{rd}$ in model size. DoPTA beats DiT-B ($94.5\rightarrow 94.6$) with less than ${1/2}$ of the pre-training steps. Our method also remains competitive with VGT despite not requiring OCR during inference, lower model size ($85$M as compared to $174$M in VGT) and lower resolution pre-training ($512$ as compared to $768$ in VGT).
@@
4.3.2 D4LA
Baselines and Results. We report the category-wise and overall mean average precision mAP@IoU[$0.50:0.95$] of bounding boxes in Table 3. We compare against DiT and VGT, the current state-of-the-art baselines. Both baselines were fine-tuned with the same setup and hyperparameters as our method. DoPTA-HR sets a new SOTA ($69.2\rightarrow 69.7$) on this benchmark, despite having less than half the parameters (85M vs 174M) and pre-training at a lower resolution (512 vs. 768) compared to VGT. In particular, DoPTA shows a marked improvement in object categories such as DocTitle, Question, ParaTitle, RegionTitle, RegionKV, Date, Author, and PageNumber. These are highly fine-grained categories, where semantic understanding of the text is crucial, highlighting the efficacy of the proposed patch-text alignment loss. We do notice a tangible performance dip ($>1\%$) in classes such as Equation, RegionList, LetterDear and LetterHead. These objects while holding individual semantic meanings could also be considered as subclasses of paragraph or list items, which might be a reason for their improper classification. Objects like equation indicate that sufficient OCR text diversity could be missing in pre-training. We present and study qualitative examples of such failure cases in Appendix C.
Expt. ID Model Patch-Text Alignment Masking Ratio RVL-CDIP PubLayNet D4LA
1 CLIP - - 90.97 93.3 64.5
2 DoPTA - 92.51 94.35 67.48
3 DoPTA 0.2 92.54 94.43 67.74
4 DoPTA 0.4 92.79 94.54 67.7
5 DoPTA 0.6 92.84 94.62 67.92
Table 4: Evaluation of performance at different masking ratios. Evaluations are done at $224\times 224$ resolution at 160k pre-training steps. Publaynet and D4LA were evaluated with default DiT config. Best results are highlighted in bold.
@@ 4.3.2 D4LA
DoPTA-HR 85M 95.29 94.18 94.73
Table 5: Text detection accuracy (IoU@0.5) on FUNSD, where Mask R-CNN is used with different backbones. Best result in each category is indicated in bold. Second best result is underlined

### v2

@@ 1 Introduction
Figure 1: Our method achieves superior FPS due to the OCR free inference setting while also setting SOTA mAP as compared to several existing methods. Model(OCR) denotes the FPS when OCR parsing is taken into account for computing inference time.
The aforementioned reasons highlight the necessity of multimodal modeling for effective document understanding. Currently, transformer architecture has evolved as a ubiquitous framework capable of modeling multiple modalities and has naturally been applied to VDU as well. Most works [48, 49, 19, 1, 43] use a unified transformer approach, wherein image, text, and layout information are processed by a single multi-modal transformer, which is pre-trained with a variety of objectives. The unified transformer based methods focus more on textual information, and treat visual information as secondary. They require extraction of text from a document image using standard OCR techniques, which is later modeled by the unified transformer for downstream tasks. This 2-stage paradigm has two key issues - inflexibility and latency of the OCR pipelines, and error propagation from OCR extraction to downstream tasks. Methods such as Donut [24] instead model both OCR extraction and subsequent understanding of the document in a single end-to-end approach using an encoder-decoder transformer model.
Although these approaches achieve strong performance on semantic tasks such as document question answering [30] and information extraction [40], they fall behind on visual tasks such as document layout analysis, as their primary focus is on modeling the textual features. In this field, state-of-the-art results have been achieved by DiT [25] and VGT [12]. DiT approaches document image understanding in a self-supervised fashion, wherein masked image patches are reconstructed through a ViT encoder, and matched to the tokens from a pre-trained dVAE. VGT builds upon this, adding a Grid Transformer (GiT) to infuse layout information into the image representations. While these methods achieve strong results, we argue that the semantic information from text in the image can be a strong factor in improving the layout understanding. Inspired by the power of contrastive language-image training in representation learning [36] and fine-grained image-text alignment techniques like FILIP [50], we specially design a patch-text alignment loss for documents, using IoU to guide the model to learn effective representations that are semantically and structurally rich. Our key contributions can be summarized as follows:
@@ 2 Related Work
Document Image Understanding. Visual document understanding(VDU) requires careful design of the objectives, owing to the unique structure of these images. The majority of approaches in this field can be categorized into two sub-categories based on the use/non-use of OCR as an input. i) OCR-Based methods include BiVLDoc [29], LayoutLM [48, 49, 19], DocFormer [1], BROS [18], VL-BERT [41], UDOP [43], VGT [12], TILT [35], M2Doc[53] and UDOC. These works utilize off-the-shelf OCR methods to parse the text and bounding boxes from a document image. The textual and image features are later combined through early or late fusion, using a joint transformer encoder to produce the final representations. Different variants of objectives such as masked image modeling (MIM), masked language modeling (MLM), and image-language alignment are proposed in these papers. However, these works require OCR as an input during inference. ii) OCR-Free methods such as Donut [24], DiT [25], and StructTextv2 [51] instead aim to learn visual features in the absence of OCR as an input during inference, though it may be utilized as a target during pre-training. Donut uses a transformer encoder-decoder architecture with OCR parsing as its pre-training task. On the other hand, DiT learns image features in the absence of any OCR ground truth, by aligning image patches with learned tokens from a discrete VAE tokenizer. StructTextv2 uses a dual objective of image reconstruction and text prediction of masked-out regions. Our work falls into the second category.
@@ 3 Methodology
We now present DoPTA, a novel pre-training method for learning document image representations with strong semantic and structural understanding. The key component of DoPTA is the introduction of a novel fine-grained image-text contrastive alignment objective for document images. This loss imbibes textual-semantic information into the image representations, leading to better structural understanding through the semantics. Despite the strong performance demonstrated by this loss, as shown in Sec. 5DoPTA also includes an image reconstruction loss to incorporate additional structural information. We present a detailed description of our losses and architecture in the following section.
@@ 3.1 Model architecture
Image Encoder. The  DoPTA encoder (EI) follows Vision Transformer [13]. We reshape an image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ into a sequence of flattened 2D patches $\mathbf{x}_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$, where $(H,W)$ is the resolution of the original image, $C$ is the number of channels, $(P,P)$ is the resolution of each image patch, and $N=HW/P^{2}$ is the resulting number of patches. We randomly mask out a fraction $M$ of the patches by replacing them with a learned [MASK] embedding, which is later reconstructed by the image decoder. Learnable positional embeddings are added to each patch before passing it as input to the transformer. The per-patch embeddings are aligned with both textual and visual information to ensure a rich representation.
Text Encoder. The text encoder (ET) is a transformer model. We extract the text with corresponding bounding boxes from each document image using an off-the-shelf OCR engine. The entire set of texts is concatenated in reading order and tokenized as a single string. We truncate the tokenized string at a maximum sequence length of $L_{T}$ tokens and add learnable positional embeddings before passing through the text encoder model. We use the per-token embeddings at the output layer for aligning vision features.
@@ 3.3 Image Reconstruction Loss
Figure 3: Heatmap visualisation of the normalised dot product similarity of image region embeddings with the text embedding for the token ‘phosphine’ taken from DoPTAmodel. Additional qualitative results are presented in Appendix B.
@@ 4 Experiments
Next, we present experimental results to show the effectiveness of image features produced by DoPTA on a variety of document tasks, including document image classification (Sec. 4.2), document layout analysis (Sec. 4.3), and text detection (Sec. 4.4). Evaluations show that our model achieve state-of-the-art results in multiple tasks, outperforming larger models, while adopting a significantly shorter pre-training schedule.
@@ 4.1 Implementation and Pre-Training
We pretrain DoPTA on the IIT-CDIP [39] dataset. This dataset contains 42M pages of black-and-white document images containing rich text. We extract word-level OCR text and their bounding boxes using EasyOCR [21] pipeline and use random cropping as the image augmentation. Though we carefully ensure the quality of the extracted OCR through filtering, some errors in extracted OCR do persist. In Appendix A, we explore the use of a PDF dataset which circumvents this issue. However, we choose the CDIP dataset for pre-training due to its larger scale, and to maintain parity with the baselines, which pre-train on the same dataset. An important point to note is that our method does not require any OCR input during inference.
We use a mix of padded (aspect ratio preserving) and square-cropped images during training to ensure good downstream performance in all settings. We follow the architectural choices of ’CLIP-ViT-B/16’, with our DOPTA encoder and text encoder being 12-layer transformer with 8 attention heads. The hidden (intermediate) sizes are 768 (3072) for the DOPTA encoder and 512 (2048) for the text encoder. Both models are initialized using the ’CLIP-ViT-B/16’ weights. The context length of the text encoder is set to $512$ by linear interpolating the learnt CLIP positional embeddings. As discussed in Sec. 3.1, we adopt a lightweight 2-layer transformer image decoder each with $8$ attention heads. We train DoPTA with image resolution $512\times 512$. The model uses a patch size of 16, a global batch size of $2048$, dropout of $0.1$, and learning rate of $31e-$ 3. The masking ratio $M$ for reconstruction is set to $0.6$. We train DoPTA  for 15 epochs ($\approx$250k steps). This is a significantly shorter pre-training schedule compared to other works like DiT [25], LayoutLMv3 [19], and VGT [12], which are pretrained for 500k steps or more.
@@ 4.2 Document Image Classification
We use the RVL-CDIP benchmark to evaluate the classification performance of DoPTA. The benchmark consists of 400K document images split into 320K train, 40K validation and 40K test images. It consists of $16$ different classes like advertisement, email, form, scientific publication, etc. To perform classification, we obtain a single representation embedding per image by average pooling the patch-wise embeddings and directly applying a linear classification head on top. We evaluate DoPTA encoder by finetuning for 100 epochs on the training set as done in DiT [25]. We use AdamW optimiser with a learning rate of $31e-$ 3, a global batch size of 1024 and perform gradient clipping with a value of $0.1$.
Baselines and Results. We compare and report results of classification accuracy on the test set in Table 1. We consider two categories of methods - i) OCR-based methods which rely on the OCR-identified text in the image as input, and ii) OCR-Free methods which treat document image classification purely in the image domain. All results are taken from the DiT[25], except Donut which we finetune ourselves. Donut [24] utilizes an encoder-decoder architecture for end-to-end OCR extraction. To test the performance of Donut in the image domain, we utilize the image encoder alone and evaluate it using the aforementioned setup. While the OCR-based methods achieve the highest performance in this category, we find that  DoPTA  outperform all OCR-free methods. It is notable that DoPTA  outperforms even the DiT-L model, despite having $<1/3^{rd}$ the parameters, and a much shorter pre-training schedule (250k steps in our case as compared to 500k steps for DiT-L).
@@ 4.3 Document Layout Analysis
For object detection, we use a Cascade R-CNN [6] as the detection pipeline on top of the backbone models, using the Detectron2 [46] library to evaluate our models. We use the same FPN and data processing setup as DiT [25] and VGT [12], with resolution-modifying modules at four different transformer blocks ($3$, $5$, $7$, and $11$) to adapt the single-scale ViT to the multi-scale FPN. Let $d$ be the total number of blocks; the $d/3^{rd}$ block is upsampled by 4×4$\times$4 × using a module with $2$ stride-two $2\times 2$ transposed convolution. For the output of the ${d/2}^{th}$ block, we use a single stride-two $2\times 2$ transposed convolution to upsample by 2×2$\times$2 ×. The output of the ${2d/3}^{th}$ block is utilized without additional operations. Finally, the output of $d^{th}$ block is downsampled by 2×2$\times$2 × with stride-two $2\times 2$ max pooling. All images are cropped with probability $0.5$ to a random rectangular patch which is then resized again such that the shortest side is at least $480$ and at most $800$ pixels while the longest is at most $1,333$ pixels.
@@ 4.3.1 PubLayNet
PubLayNet [54] is a large dataset of 360K images for document layout analysis, created from over one million scientific articles in PubMed Central. It includes labeled images with five layout regions: text, title, list, figure, and table. We finetune our model on the training split $(335,703)$ and evaluate on the validation split $(11,245)$. We follow the setting of DiT [25] and train for 60K steps with a batch size of $16$ and a learning rate of $44e-$ 4.
Baselines and Results. We report the category-wise and overall mean average precision mAP@IoU$[0.50:0.95]$ of bounding boxes in Table 2. We compare with vision-only input models such as DiT and Donut, as well as vision+OCR input models like VGT, LayoutLMv3. DoPTA  achieves 94.9 which is on-par with DiT-L despite being less than ${1/3}^{rd}$ in model size and with less than ${1/2}$ of the pre-training steps. Our method also remains competitive with VGT despite not requiring OCR during inference, lower model size ($85$M as compared to $174$M in VGT). From our observations, classes like text, list and table are more ambiguous on textual semantics and might not be the best place to test our patch-text alignment loss. We see better performance gains on the D4LA and M6Doc benchmarks which has classes with more semantic distinction in the following sections.
@@
DoPTA 85M 95.29 94.18 94.73
Table 4: Text detection accuracy (IoU@0.5) on FUNSD, where Mask R-CNN is used with different backbones. Best result in each category is indicated in bold
4.3.2 D4LA
Baselines and Results. We report the category-wise and overall mean average precision mAP@IoU[$0.50:0.95$] of bounding boxes in Table 7. We compare against DiT and VGT, the current state-of-the-art baselines. Both baselines were fine-tuned with the same setup and hyperparameters as our method.  DoPTA  sets a new SOTA ($69.2\rightarrow 70.72$) on this benchmark, despite having less than half the parameters (85M vs 174M) and pre-training at a lower resolution (512 vs. 768) compared to VGT. In particular, DoPTA shows a marked improvement in object categories such as DocTitle, Question, ParaTitle, RegionTitle, RegionKV, Date, Author, and PageNumber. These are highly fine-grained categories, where semantic understanding of the text is crucial, highlighting the efficacy of the proposed patch-text alignment loss. We do notice a tangible performance dip ($>1\%$) in classes such as Equation, RegionList, LetterDear and LetterHead. These objects while holding individual semantic meanings could also be considered as subclasses of paragraph or list items, which might be a reason for their improper classification. We also observe a significant class imbalance for objects like equation which results in a wide variation in performance. Deeper analysis on the predictions of DoPTA  on equation class revealed that  DoPTA  was able to detect chemical equations which were originally not present in ground truth and was not predicted by VGT. We present and study qualitative examples of such cases in Appendix C.
Model Patch-Text Masking RVL-CDIP PubLayNet D4LA
Alignment Ratio
CLIP - - 90.97 93.3 64.5
DoPTA - 0.6 92.3 94.35 66.7
DoPTA - 92.51 94.35 67.48
DoPTA 0.6 92.84 94.62 67.92
Table 5: Evaluation of performance with loss combinations. Evaluations are done at $224$ resolution at 160k pre-training steps. Publaynet and D4LA were evaluated with default DiT config. Best results are highlighted in bold.
Masking Ratio RVL-CDIP PubLayNet D4LA
0.2 92.54 94.43 67.74
0.4 92.79 94.54 67.7
0.6 92.84 94.62 67.92
Table 6: Evaluation of performance at different masking ratios at 160k pre-training steps. Best results are highlighted in bold.
Method AP50 AP75 mAP
Mask R-CNN 58.4 46.2 40.1
Cascade R-CNN 70.5 62.9 54.4
HTC 74.3 67.2 58.2
SCNet 73.5 65.1 56.1
Deformable DETR 76.8 63.4 57.2
ISTR 80.8 70.8 62.7
TransDLANet 82.7 72.7 64.5
VSR 76.2 68.8 59.9
DINO 84.6 76.7 68.0
M2Doc 78.0 70.7 61.8
DiT-B 84 76.4 67.6
DoPTA 85.8 78.5 69.5
Table 7: Performance comparison of different methods on M6Doc benchmark. Uses OCR as input during inference.Best result in each category is indicated in bold
4.3.3 M6Doc
M6Doc is another layout detection benchmark with highly nuanced object classes like poem, examinee information, weather forecast that could rely on semantic understanding for efficient detection. The dataset constitutes of $9080$ images with $76$ highly diverse object classes. There are also objects like QR Code, flag, underscore which are more visually distinguishable. We finetune our model for $90$K steps with a batch size of 16 and a learning rate of $44e-$ 4 with a warmup of $100$ steps using the Cascade-RCNN framework.
Baselines and Results. We report the AP50, AP75 and mAP scores of our model and baselines on the validation set. Except DiT-B, all the baselines are reported from M2Doc[53]. We train DiT-B with the same hyperparameters as ours.  DoPTA beats DiT-B by +1.9 mAP DoPTA achieves $69.5$ mAP setting SOTA on the benchmark.